A critical input to all stock valuation models is the discount rate. Group of answer choices

True

False
Answer #1

Discount rate is a critical input to all stock valuation models - True

When we discount a particular cash flow or a value, it has the compounding effect on the valuation. Thus even a small change in the discount rate has a large effect on the valuation

As in the example below, a change of just 1% can change the valuation end result

1 2 3 4 5 Terminal Value N Discount rate Discount rate 10000 10 8.0% 9.0% 6 7 at 8% at 9% 4632 =B2/(1+B4)^B3 4224 =B2/(1+B5)^